# Grounding Synthetic Data Evaluations of Language Models in Unsupervised Document Corpora

### v1

@@ Abstract
Language Models (LMs) continue to advance, improving response quality and coherence.Given Internet-scale training datasets, LMs have likely encountered much of what users might ask them to generate in some form during their training. A plethora of evaluation benchmarks have been constructed to assess model quality, response appropriateness, and reasoning capabilities. However, the human effort required for benchmark construction is limited and being rapidly outpaced by the size and scope of the models under evaluation. Additionally, having humans build a benchmark for every possible domain of interest is impractical. Therefore, we propose a methodology for automating the construction of fact-based synthetic data model evaluations grounded in document populations. This work leverages those very same LMs to evaluate domain-specific knowledge automatically, using only grounding documents (e.g., a textbook) as input. This synthetic data benchmarking approach corresponds well with human curated questions with a Spearman ranking correlation of $0.96$ and a benchmark evaluation Pearson accuracy correlation of $0.79$. This novel tool supports generating both multiple choice and open-ended synthetic data questions to gain diagnostic insight of LM capability. We apply this methodology to evaluate model performance on a recent relevant arXiv preprint, discovering a surprisingly strong performance from Gemma3 models.
@@ 1 Introduction
Recent advances in language models (LMs) have demonstrated significant capability improvements in natural language processing, generation, and understanding across diverse domains. Frontier models demonstrate remarkable proficiency over a variety of tasks, from semantic text summarization to creative ideation conditioned upon input prompts. This expansion in capability has been enabled primarily by training on Internet-scale datasets, which expose these models to vast quantities of raw text (of varying quality levels) that summarize and compress a significant portion of human discourse. The usefulness of these AI tools rests upon their ability to generate plausible text. However, this is a double edged sword, as what sounds plausible might not be true, or even grounded in the facts of a situation. As AI tools improve workflows in professional settings, it is imperative to understand the domain specific model capability.
The evaluation of LMs presents a significant challenge. Traditional multiple choice question (MCQ) benchmarks such as MMLU [1] have provided valuable insights into model performance, but are static, public, and limited in scope compared to the models under evaluation. The fact that these benchmarks are static and public —and therefore can appear in training data— is a significant weakness. Combined with the human effort required to construct new, high quality tests [2], it is unlikely the evaluation construction can keep pace with LM development. This scaling challenge is particularly acute when evaluating domain-specific knowledge for professional applications. While general knowledge benchmarks are well-established, assessing model performance outside of those knowledge bases remains difficult. Understanding what elements of a given specialty domain the model can produce reliable output for and what topics it will under-perform on will accelerate high-quality professional adoption. Unfortunately, the creation of expert-validated benchmarks for every conceivable domain is neither practical nor sustainable. Therefore, it is necessary to automate the creation of high quality benchmarks grounded in facts and document corpora. Such methods will need to generate test datasets which are comprehensive with respect to the grounding documents (textbooks/manuals) while being accurate, with minimal to zero human intervention beyond document selection.
This work presents a novel approach to automated benchmark generation which addresses these challenges. By starting from high quality, authoritative grounding documents, we can ground synthetic data evaluations in facts and domain knowledge, while providing a proxy to a human curated evaluation for the topic covered by the source documents. This approach reduces human effort (especially in tedious data annotation tasks) and enables rapid adaptation to new domains via introduction of appropriate grounding documents. This project is organized around the following questions:
@@
    How can we leverage LMs to extract relevant information and generate multiple choice and open ended questions about knowledge claims from the grounding documents?
@@
    How to evaluate whether this automatic synthetic evaluation approach reproduces benchmarks in domains for which human curated evaluations exist?
@@
    What pitfalls require mitigation for successful application of synthetic data to LM capability evaluation?
@@ 2 Related Works
This approach draws inspiration from NP-completeness, where easy to verify solutions might be hard to generate. The LM under evaluation has a harder job than the LM converting existing documents into appropriately formatted test questions with accompanying answers. The evaluation LM must produce the correct information from just the question; by contrast, the benchmark creation LM has the source material. In other words, given the grounding paragraph, a language model should be much more capable at converting that textual information into a multiple choice or short answer question than it would be at zero-shot answering the very same question. This task difficulty discrepancy provides hope that this approach will produce viable results long term. Asking an LM to generate diagnostically useful evaluation benchmarks drawn only from information stored in model weights is difficult, but conversion of information from a high quality source into test questions is attainable.
This work draws on the rich history of language model evaluation, where LMs are already being used to judge and score the results of human curated benchmarks. When scoring outputs of an LM producing a multiple choice answer, it’s often useful to have another LM reason about and extract the specified [A,B,C,D] value, instead of hoping the model under evaluation only produced the exact match answer letter [3, 4, 5]. This process becomes even more important when evaluating short answers from LMs instead of multiple choice responses, or when scoring the responses and providing reasoning as to why an answer was correct or incorrect [6, 7, 8]. This idea of automating question answer generation has been approached before. The most similar contemporary work is YourBench by Shashidhar et al. [9], which leverages the convergent idea of grounding documents supporting LM benchmark generation; demonstrating the synthetic replication of MMLU. Additionally, AutoBencher by Li et al. explores leveraging synthetic data (supported by privileged information) for model evaluation with RL to optimize the generated evaluation for discovering gaps in LM capability [10]. The use of grounding documents or privileged information greatly improves the usefulness of the synthetic data benchmark, as you are no longer simply asking another LM for questions (and answers).
Early Question Answer Generation (QAG) approaches attempted to solve the question creation from a context + correct answer input. These approaches include improvements in selecting better LM samples [11], increasing sample diversity over distributional shifts [12], explorations of interrogative prompting against the context [13], and evaluating round trip information consistency [14]. These early methods sometimes fine-tune generation models to improve task performance, though largely that is only done with smaller models [15, 16]. Some work focused on fine-tuning specialty models to improve option generation for multiple-choice questions [16]. Other early work during the shift to general purpose LMs involved leveraging those language models directly as data generation sources. This took the form of prompting the model to write questions (without grounding context) [17, 18, 19, 20], or configuring an LM-as-evaluator/interrogator dynamically interacting with the LM under evaluation to explore capability in a reference-free manner [17]. All of these approaches attempt to address the dataset leakage problem where benchmarks, once published, become part of the training dataset for the next round of models. As the capability of LMs increases over time, setting up adversarial LM-evaluator vs LM-evaluatee systems is becoming more feasible [21]. Though extending those systems such that the LM-evaluator has grounding documents to pull from should strengthen the performance.
Scaffolding Question Generation with Grounding Documents.
With the rise of general purpose token-tumblers like ChatGPT, longer, more complex grounding documents become increasingly useful. There is no longer a need to constrain synthetic benchmark creation to a short paragraph context + correct answer input; the LM can operate on larger, more complex document chunks, with less pre-processing required to make the input context data useful. Contemporaneous work YourBench [9] and AutoBencher [10] can leverage lightly-filtered web-pages as the context, chunking the data into rough paragraphs for ingestion into the question generation and grounding system. The SecQA benchmark was created partly through synthetic data grounded in an open-source computer security textbook [22]. Similarly (though much earlier), the PubMedQA benchmark was created through human and synthetic data generation grounded in biomedical research paper abstracts [23]. Some research demonstrates that operating purely on document summaries produces inferior generated questions compared to leveraging the full document [24]. With modern long-context LMs, summarization is becoming less of a problem, as models can manipulate significant volumes of text within their attention context. Synthetic data question generation systems have seen rapid improvement. This is partly driven by capability improvements in the LMs powering synthetic data generation. However, there has also been exploration into improvements in prompting for diverse question generation [25] and breadth of topic generation for chatbot arena style questions [26]. Enhancing retrieval augmented generation (RAG) systems for domain specific applications has also used synthetic data question generation to compare baseline RAG to fine-tuning [27].
LM as Feature Extractors.
There are many aspects of text that one might want to quantify, however imprecisely. However, devoting the required copious human annotation hours is just impractical. Current LMs end up being acceptable feature extractors, when provided detailed prompts. Chiang and Lee demonstrate that LMs can substitute for human evaluators in assessing text quality with strong correlation in the results [28]. Similarly, Han et al. leverage LM evaluators for judging quality, finding the results highly correlate with human judgments [29]. The use of LMs to extract features from text is not to dissimilar from using the LM-as-a-Judge pattern for large scale open ended capability evaluations [30, 31, 32]. The LM is capable of grading open ended questions for correctness, so determining features other than correctness form a text sample is viable. However, there are limits on what LM-as-a-Judge can is capable of, especially in the field of information retrieval [33, 32]. Relying on LMs to extract properties like answer correctness given sufficient scaffolding and context is a well-used pattern.
Multiple Choice vs Open Ended.
When constructing knowledge or fact based evaluation benchmarks, there are two broad categories of answers: multiple choice, and open ended. Open-ended questions provide the model less information and require synthesis [34], but at the cost of needing another LM to evaluate the correctness of the results (LM-as-a-Judge [30][35]. Multiple choice question (MCQ) results can be parsed and graded more simply. Modern evaluation frameworks like Inspect-AI [36] enable shuffling answer options and contain default prompts for both open-ended and MCQ evaluations. With MCQ, having the answers visible to the LM can artificially inflate performance [37], even going as far as the model identifying the correct answers without seeing the question [38]. Both open-ended and multiple choice questions have weaknesses to question reformatting and phrasing where semantically identical questions produce different answers [39]. The simplest form of open-ended evaluation relies on a judge LM, though other frameworks have been proposed which attempt to improve the precision of open-ended evaluations [40, 41].
@@ 3 Methods
Constructing synthetic data benchmarks from domain-specific relevant documents requires document chunking into contexts, topic extraction per context, question and answer generation, and finally an evaluation framework to determine model capability on the newly minted topic grounded benchmark.
@@
Modern LMs have long context windows which could support whole document ingestion, however chunking encourages the generated questions to be spread throughout the grounding document. An input document is first converted into markdown for interoperability with LMs. We use Docling [42] for conversion from pdf to markdown, and pandoc [43] for conversion from html to markdown. Once in markdown, each document section is used as a context chunk for question generation. This chunking method is simple and relies on the document author crafting relevant sections. Improvements to chunking and grounding document data preparation will likely produce overall system improvements, but were not required to demonstrate baseline performance in this study. For instance, our approach only leverages the textual data; formatting and including richer sources of information from table, charts, and images could improve evaluation quality.
@@
For each extracted document context chunk, an LM is used to extract and propose any topics present which would make high quality difficult questions for evaluating domain expert knowledge about the information in context chunks. This topic extraction leverages the same LM as the later question generation process. We explored using an embedding model to cluster topics and de-duplicate, but empirically found no satisfying improvements, as the topic generation by LM does a reasonably good job of de-duplication. From manual observations, the only duplicate topics would stem from different source document context chunks. This duplication is acceptable as the question generation is grounded in the context chunk, so any duplicate topics across context chunks will produce different questions.
@@
Given (a) the document context chunk and (b) a single topic for a question, the generating LM is prompted to construct a high quality and difficult evaluation question to probe information understanding of an expert in the domain. During question generation the LM is required to write the question, correct answer, and a justification/explanation for why the correct answer is in fact right. The answer explanations are not used in the evaluation benchmark, and simply exist to guard against models generating incorrect answers. In summary, the single grounding document is broken into chunks of context, then per-context relevant topics are extracted, then per-context/topic pair, a single evaluation question is created. This results in a list of questions with correct answers grounded in (and derived from) the source document. An example drawn from the Squadv2 dataset [44] shows a question generated from only the context by Llama-3.3-70B-Instruct:
@@
In this study we explored both multiple choice and open-ended question generation, each requiring slightly differing LM prompts. While simply using the correct multiple choice answer option in an open-ended configuration would work, we observed open-ended questions enabling more variety from the generating model if prompts are modified. One study goal was to determine if any differences manifested between multiple choice and open-ended questions for synthetic data model benchmarking. The difference in response grading requirements between open and multiple-choice could favor MCQ for fast and simple exploration of model weaknesses.
@@ 3.1 Benchmark Evaluation
Given a list of questions and answers, an LM evaluation framework (Inspect-AI [36]) is used to determine performance for the models under evaluation. The benchmark consists of many questions with associated answers. The resulting benchmark evaluation accuracy is the average number of correct answers divided by the total number of questions. LMs are evaluated against both the synthetic evaluation benchmark dataset and the original questions. The overall accuracy per evaluation model can be compared for both the synthetic and original questions across multiple datasets of grounding documents. The text of the various benchmark questions and answers are never directly compared, only the resulting LM performance on the benchmark evaluations. Figure 2 demonstrates overall system performance of synthetic multiple choice questions averaged across datasets and generating models, with a Spearman ranking correlation of $0.961$, while demonstrating slight model over-performance on the synthetic benchmark.
Figure 2: Comparison of model performance across datasets with the multiple choice question synthetic data benchmarks created by ensembling the top 5 most capable LMs. X-axis is sorted by the average values per model.
@@ 4 Evaluation Methodology
Evaluation of the quality and utility of the newly minted synthetic data benchmark built from grounding documents is done by comparing the synthetic data pipeline on datasets where human annotated questions and correct answers exist for known context paragraphs. This provides a directly comparable result, both for the overall benchmark accuracy (between human and grounded synthetic data) and the rank ordering of those models the benchmark is evaluated on. We leverage five Natural Language Processing (NLP) datasets where each entry contains a context paragraph, human written question, and correct answer. For each dataset a variety of LMs are used to generate the grounded synthetic benchmark. Per-question quality features are extracted and holistic benchmark-level question diversity is measured and compared to the human written questions. Finally, an automatic quality control pass is completed to grade both question and answer correctness/validity, in order to investigate and characterize which benchmark-generating models can write correct/valid question and answer tuples.
@@ 4.1 Datasets
Evaluating the validity of generated synthetic data questions requires source documents which have human annotated question/answer pairs to compare against. The following NLP datasets were selected:
  1. 1.
    Squadv2 [44]
  2. 2.
    HotpotQA [45]
  3. 3.
    NaturalQuestionsShort [46]
  4. 4.
    PubMedQA [23]
  5. 5.
    SecQA [22]
These datasets collectively contain of tens of thousands of context, question, answer tuples. Accordingly, a random subset of 1000 tuples per dataset was sampled as the grounding context document. For certain datasets, pre-processing was required to clarify, curate, and reformat the source human written questions into self-contained questions. For some questions this required disambiguation so the question is answerable without the context. For the multiple choice evaluation, three believable incorrect answer options needed to be created. The PubMedQA [23] dataset required conversion from yes/no/maybe questions into multiple-choice or open-ended. Only the SecQA [22] dataset already contained all of the required information and was not pre-processed. All other datasets went through an LM based reformatting operation to disambiguate the questions and (for multiple choice) invent additional incorrect answers. Table 1 provides an example of the reformatting operation for a few questions related to a single context within the Squadv2 dataset. The context is unmodified from the source dataset, and the original question and answer are shown next to the reformatted/clarified question with the associated answer choices.
@@
What is the AFC short for? American Football Conference What does AFC stand for in the context of the NFL? (A) American Football Championship (B) American Football Conference (C) American Football Council (D) American Football Committee
@@
Table 1: Example reformatted questions (with three incorrect answer choice distractors) drawn from SQuADv2.
Three different models (Llama-4-Maverick-17B-128E-Instruct, gemma-3-27b-it, o4-mini) were used for this reformatting operation and manual inspection of the results showed minimal impact related to which model was used. The remainder of this paper shows results based on the Llama-4-Maverick-17B-128E-Instruct reformatted questions.

### v2

@@ Abstract
Language Models (LMs) continue to advance, improving response quality and coherence. Given Internet-scale training datasets, LMs have likely encountered much of what users may ask them to generate in some form during their training. A plethora of evaluation benchmarks have been constructed to assess model quality, response appropriateness, and reasoning capabilities. However, the human effort required for benchmark construction is rapidly being outpaced by the size and scope of the models under evaluation. Having humans build a benchmark for every possible domain of interest is impractical. Therefore, we propose a methodology for automating the construction of fact-based synthetic data model evaluations grounded in document populations. This work leverages the same LMs to evaluate domain-specific knowledge automatically, using only grounding documents (e.g., a textbook) as input. This synthetic data benchmarking approach corresponds well with human curated questions producing a Spearman ranking correlation of $0.97$ and a benchmark evaluation Pearson accuracy correlation of $0.75$. This novel approach supports generating both multiple choice and open-ended synthetic data questions to gain diagnostic insight of LM capability. We apply this methodology to evaluate model performance on two recent arXiv preprints, discovering a surprisingly strong performance from Gemma-3 models on open-ended questions. Code is available at https://github.com/mmajurski/grounded-synth-lm-benchmark
@@ 1 Introduction
Recent advances in language models (LMs) have demonstrated significant capability improvements in diverse natural language processing tasks. This expansion in capability has primarily been enabled by training on Internet-scale corpora. As AI tooling improves professional workflows, rigorous evaluation is imperative to understand the domain-specific model capability.
Traditional multiple choice question (MCQ) benchmarks such as MMLU [1] offer valuable general insights into model performance, but are static, public, and limited in scope. The fact that these benchmarks are static and public—and therefore can appear in training data—is a significant weakness. Combined with the human effort required to construct new high quality tests [2], it is unlikely evaluation construction can keep pace with LM development. This scaling challenge is especially acute for domain-specific professional applications, where the general benchmarks provide limited insight. Unfortunately, creating expert-validated benchmarks for every domain is impractical. New automated methods must generate accurate, comprehensive, diagnostically relevant evaluations grounded in authoritative corpora with minimal human intervention beyond document selection.
We propose an novel automated methodology for generating synthetic benchmarks grounded in authoritative documents. This approach reduces human effort and enables rapid adaptation to new domains by generating fact-based multiple choice or open-ended questions directly from appropriate grounding documents. This work is organized around three core questions:
@@
    How can LMs extract relevant information and generate high-quality questions grounded in domain documents?
@@
    Do these synthetic benchmarks replicate existing human-curated evaluations?
@@
    What limitations require mitigation for successful synthetic data evaluation of LM capability?
@@ 2 Related works
Our approach is motivated by task asymmetry similar to NP-completeness: given grounding documents, an LM should reliably generate diagnostically relevant questions, while zero-shot answering the same questions is substantially harder. This information asymmetry motivates using LMs to generate benchmarks grounded in curated documents rather than relying only on model weight knowledge. Language models are routinely adopted as judges: extracting answer choices [3, 4], grading open ended responses [5, 6], and justifying correctness [7]. The most similar contemporary convergent work is YourBench [8] synthetic replication of MMLU from grounding documents and AutoBencher [9] reinforcement learning optimization to discover gaps in LM knowledge. Leveraging grounding documents greatly improves synthetic data utility, as the methodology no longer relies on internalized LM knowledge.
Modern long-context LMs can process entire documents. Benchmarks like SecQA [10] and PubMedQA [11] leverage domain-specific textbooks and web pages. Full-document grounding generally outperforms summary-based systems [12]. Continued LM improvement, prompting strategies [13], and topic breadth [14] has further improved synthetic data. These methods also enhance downstream applications like retrieval-augmented generation [15].
LMs can serve as acceptable feature extractors when guided by detailed prompts. Studies show strong correlation between LM-based and human judgments of text quality [16, 17]. This aligns with the LM-as-a-Judge paradigm used for open-ended benchmark evaluation [18, 19]. Beyond open ended answer correctness, LMs can extract other textual features. However, limitations in LM ground truthing remain, especially in tasks like information retrieval where relevance judgments are more complex [20, 21]. LM evaluation benchmarks generally use either multiple choice (MC) or open-ended (OE) formats. Open-ended questions require synthesis [22] but depend on a judge LM for scoring [18, 23], while MCQs are easier to grade but may inflate performance due to answer visibility [24, 25]. Frameworks like Inspect-AI [26] support both types with prompt templates and answer shuffling. Both formats are sensitive to rewording, leading to inconsistent outputs [27], and newer frameworks aim to improve open-ended scoring precision [28, 29].
@@ 3 Methods
Constructing synthetic data benchmarks from relevant domain-specific documents requires chunking into contexts, topic extraction per context, question and answer generation, and finally an evaluation framework to determine model capability on the newly minted benchmark. All LM interactions use default generation parameters, except we set $\text{temperature}=0.7$ (o4-mini did not support setting temperature).
@@
While modern LMs have sufficient context windows to support whole document ingestion, chunking encourages the generated questions to be spread throughout the grounding document. An input document is first converted into markdown for interoperability with LMs. We use Docling [30] for conversion from pdf to markdown, and pandoc [31] for conversion from html to markdown. Once in markdown, each document section is used as a context chunk for question generation. This chunking method is simple and relies on the document author crafting relevant sections. Enhancements to chunking and document data preparation will likely produce overall system improvements, but were not required to demonstrate baseline performance in this study. For instance, our approach only leverages the textual data; formatting and including richer sources of information from table, charts, and images may improve evaluation quality.
@@
For each extracted document context chunk, an LM is used to extract and propose any topics present which make high quality difficult questions for evaluating domain expert knowledge about the information in context chunks. This topic extraction leverages the same LM as the question generation process. We explored using an embedding model to cluster topics and de-duplicate, but empirically found no satisfying improvements, as the topic generation by LM does a reasonably good job of de-duplication. From manual observations, the only duplicate topics stem from different source document context chunks. This duplication is acceptable as the question generation is grounded in the context chunk; any duplicate topics across context chunks will produce different questions.
@@
Given (a) the document context chunk and (b) a single topic for a question, the generating LM is prompted to construct a high quality and difficult evaluation question to probe the knowledge and understanding of an expert in the domain. During question generation, the LM is required to write the question, correct answer, and a justification/explanation for why the correct answer is in fact right. The answer justifications are not used in the evaluation benchmark and simply exist to guard against generating incorrect answers. In summary, the single grounding document is broken into chunks of context, per-context relevant topics are extracted, then per-context/topic pair, and a single evaluation question is created. This results in a list of questions with correct answers grounded in (and derived from) the source document. This example drawn from Squadv2 [32] shows a question generated by Llama-3.3-70B-Instruct grounded in the context:
@@
In this study, we explored both multiple choice and open-ended question generation, each requiring slightly differing LM prompts. While simply using the correct multiple choice answer option in an open-ended configuration works, we observed that open-ended questions enable more variety from the generating model if prompts are modified. One study goal was to determine if any differences manifested between multiple choice and open-ended questions for synthetic data benchmarking. The difference in response grading requirements between open and multiple-choice favors MCQs for fast and simple exploration of model weaknesses.
@@ 3.1 Benchmark evaluation
Given a list of questions and answers, an LM evaluation framework (Inspect-AI [26]) is used to determine performance for the models under evaluation. The benchmark consists of many questions with associated answers. The resulting benchmark evaluation accuracy is the average number of correct answers divided by the total number of questions. LMs are evaluated against both the synthetic evaluation benchmark and the original questions. The overall accuracy per evaluated model can be compared for both the synthetic and original questions across multiple datasets of grounding documents. The text of the various benchmark questions and answers is never directly compared, only the resulting LM performance. Figure 2 showcases the performance of synthetic multiple choice questions averaged across datasets and generating models, with a Spearman ranking correlation of $0.968$, demonstrating slight model over-performance on the synthetic benchmark.
Figure 2: Comparison of average LM performance across six datasets with the multiple choice question synthetic data benchmarks created by the ensemble of bolded model in Table 2.
@@ 4 Evaluation methodology
Evaluation of the quality and utility of the newly minted synthetic data benchmark built from grounding documents is done by comparing the synthetic data pipeline on datasets where human annotated questions and correct answers exist for known context paragraphs. This provides a directly comparable result, both for the overall benchmark accuracy (between human and grounded synthetic data) and for the rank ordering of those models that the benchmark is evaluated on. We leverage six Natural Language Processing (NLP) datasets where each entry contains a context paragraph, human written question, and correct answer. For each dataset a variety of LMs are used to generate the synthetic benchmark. Per-question quality features and benchmark-level question diversity is extracted to compare against the human written questions. Finally, an automatic quality control pass is completed to grade both question and answer correctness/validity, in order to investigate and characterize which benchmark-generating models can write valid question and answer pairs.
@@ 4.1 Datasets
Evaluating the validity of generated synthetic questions requires source documents which have human annotated question/answer pairs to compare against. The following six NLP datasets were selected: (a) Squadv2 [32] (b) HotpotQA [33] (c) TrivaQA-web [34] (d) NaturalQuestionsShort [35] (e) PubMedQA [11], and (f) SecQA [10].
Table 1: Example human question reformatting
@@
These datasets collectively contain tens of thousands of context, question, answer tuples. A random subset of 100 tuples per dataset was sampled as the grounding context document. For certain datasets, pre-processing was required to clarify, curate, and reformat the source human written questions into self-contained questions. For some questions this required disambiguation so the question is answerable without the context. For the multiple choice evaluation, three believable incorrect answer options needed to be created. The PubMedQA [11] dataset required conversion from yes/no/maybe questions into multiple-choice or open-ended. Only the SecQA [10] dataset already contained all of the required information and was not pre-processed. All other datasets went through an LM based reformatting operation correct these deficiencies. Table 1 provides an example of the reformatting operation for a few questions drawn from a single context of the Squadv2 dataset.
Three different models (Llama-4-Maverick-17B-128E-Instruct, gemma-3-27b-it, o4-mini) were used for this reformatting operation. The remainder of this paper shows results based on the Llama-4-Maverick-17B-128E-Instruct reformatting (see Table 4 in Appendix E for results of all three reformatting models).